I have a homework problem that I am not sure I am answering correctly
750 workers were surveyed about where they had attended school. 480 of the workers surveyed had attended a community college. 420 of the workers surveyed had attended a 4 year school. 75 of the workers surveyed had attended neither.
a) find the probability that a student attended a community college or attended a university
b) find the probability that a student did not attend a community college or did not attend a university.
1
Expert's answer
2012-05-18T11:38:02-0400
a) find the probability that a student attended a community college
or
attended a university
The number of workers attended a community
college or attended a university:
750 - 75 = 675
So, the probability is P
= 675 / 750 = 90%
b) find the probability that a student did not attend a
community
college or did not attend a university.
The number of workers
attended either a community college and a university:
420 - x + 480 =
675
x = 225
So, the number of workers did not attended a community college
or did
not attend a university is 750 - 225 = 475
And the probability is P
= 475 / 750 = approx. 63%